# Accounting for motion of supernova host galaxy in statistical inference from SNIa data

### v1

@@ 2.1.1 Estimator for nonlinear models
Regularization is commonly used in fitting complex models with a large number of parameters as it acts as a constraint and facilitates stable maximization of the posterior distribution (e.g. [26]). However, as we note later, the addition or omission of this term may not always change the dimensionality of the fit, as explained further in the context of the specific cosmological application of this method presented in the paper.
3 Magnitude–Redshift relation
@@
Although the measurement error in the observed redshift of a supernova host galaxy is negligibly small, due to the possible peculiar motion of the host galaxy with respect to an observer at rest with respect to the universe, their observed redshift is shifted with respect to the theoretical cosmological redshift that encodes the expansion history of the universe. To estimate this shift, we note that
$\frac{\nu_{0}}{\nu_{\rm obs}}=\frac{\nu_{0}}{\nu_{\rm com}}\frac{\nu_{\rm com}}{\nu_{\rm obs}}\approx(1+z)(1+{v_{p}}/{c})\approx 1+z+{v_{p}}/{c}$ (3.1)
@@
where $\nu_{\rm obs},\nu_{0}$, and $\nu_{\rm com}$ are the frequencies observed at earth, in the rest frame of the galaxy and a comoving observer at the position of the galaxy, and $v_{p}$ is the component of velocity of a galaxy with respect to the comoving observer parallel to the line-of-sight from the observer to the supernova host galaxy. The approximation in the last equation assumes $v_{p}/c\ll 1$, so we can neglect the transverse Doppler effect. For small redshifts where this correction dominates, it suffices to ignore the $zv_{p}/c$ term, leading to the final form.
@@
Peculiar motions also cause beaming of light towards the observer, alter the observed flux, and cause additional k-correction (e.g. see [27] for the combined impact of all these effects on the number count). The first two effects act in concert, as an object moving towards an observer would appear closer and less redshifted. The standard analysis already takes into account measurement errors in the luminosity of an SNIa, and these errors are much larger than what we expect from the peculiar motion of host galaxies. However, this is not true for redshift measurement errors that, as already noted, are small compared to peculiar motions. At low redshifts, this effect is enhanced as peculiar redshift/blueshift can become comparable to the redshift of the host galaxy. As more and more supernovae are observed, the effective statistical error in luminosity distance reduces at any redshift due to statistical averaging. Thus, we need techniques that allow us to handle errors on redshift, even if they are small.
@@
$m=M+5\log\left(\frac{D_{\rm L}}{\text{Mpc}}\right)+25\;.$ (3.2)
@@
$D_{\rm L}=\frac{c(1+z)}{H_{0}}\int_{0}^{z}\frac{dz^{\prime}}{h(z^{\prime})},$ (3.3)
@@
$h(z)=\left[\Omega_{\rm M}(1+z)^{3}+\Omega_{\rm DE}(1+z)^{3(1+w)}\right]^{1/2}\,,$ (3.4)
where $\Omega_{\rm M}$ and $\Omega_{\rm DE}$ are the present-day matter and dark energy densities respectively. The dark energy and pressure are related through the parameter $w$ through $p_{\rm DE}=w\rho_{\rm DE}$. If dark energy is assumed to be the cosmological constant, then $w=-1$ in this parametrization. The apparent magnitude can be written as
$m=M+5\log(cH^{-1}_{0}/{\rm Mpc})+5\log\left[(1+z)\int_{0}^{z}\frac{dz^{\prime}}{h(z)}\right]+25\,.$ (3.5)
@@
4 Results
In cosmological studies, our primary concern is with the parameters $w$ and $\Omega_{\rm M}$. For presenting our results, we need to marginalize various parameters, which means integrating them in the posterior distribution. Also note that in Eq. (2.11), we need to integrate over $x^{*}$, which corresponds to the redshift $z$ in this case. Clearly, it is not feasible to integrate over so many variables. Therefore, to carry out marginalization, we employ the Markov Chain Monte Carlo (MCMC) sampling technique. Specifically, we develop and implement the Metropolis-Hastings algorithm within the MCMC framework to draw samples from the posterior distribution and use GetDist [28] python library for the analysis of MCMC samples.
Figure 1: The figure displays one-dimensional posterior distributions and two-dimensional (1- and 2-$\sigma$) contours of $\Lambda$CDM parameters from Pantheon data using three different estimators. Estimator $\mathcal{E}_{1}$ (red) ignores any contribution to redshift from the peculiar motion of supernovae host galaxies, $\mathcal{E}_{2}$ considers it for the linearized model (see Appendix A for details), while $\mathcal{E}_{3}$ (green) represents our estimator, which incorporates the contribution from peculiar motion for the exact model.
Parameters Priors Best-fit values
@@ 4 Results
$\Omega_{\rm M}$ $\mathcal{U}(0.2,0.4)$ $0.286\pm 0.012$ $0.296\pm 0.014$ $0.302\pm 0.016$
$H_{0}$ $\mathcal{U}(60,80)$ $73.59\pm 0.229$ $73.34\pm 0.287$ $73.13\pm 0.343$
@@
Although distant SNIa are not ideal for determining $H_{0}$ due to degeneracy with $M$, in the following sections, we also present results for the Hubble parameter for the $\Lambda$CDM model. In order to break the aforementioned degeneracy, we use the absolute magnitude $M=-19.2478$ corresponding to the SH0ES [29] value of $H_{0}=73.04$ km/s/Mpc for the Pantheon sample.
For the redshift error, we assume that each observed $z_{i}$ in the data deviates from its true value $z^{*}_{i}$, which falls within a specified uncertainty range $z^{*}_{i}\in[z_{i}^{\rm min},z_{i}^{\rm max}]$, where the min and max values are chosen to correspond to our assumptions about peculiar velocities. This is imposed even if we assume the distribution of redshift to be nonuniform. From analytical theory or numerical simulations, we can estimate the variance of galaxy peculiar velocities as a function of redshift. Also, note that only the line-of-sight component of the peculiar velocity contributes to the change in redshift. By the isotropy of the peculiar velocity dispersion, we have $\sigma_{v_{\rm 3D}}=\sqrt{3}\sigma_{v_{p}}$
In our analysis below, we consider three distinct cases to present our results. The standard case is labeled as $\mathcal{E}_{1}$, where we do not account for the redshift error. In Appendix A, we derive an estimator based on the linearized magnitude–redshift relation labeled $\mathcal{E}_{2}$ to include the redshift error using the likelihood function (Eq. (A.6)). This is of interest since the MCMC-based final estimator ($\mathcal{E}_{3}$) using the method described in Section 2.1.1 (Eqs. (2.2)–(2.11)), which includes the contribution of peculiar velocity to the observed redshift by treating each redshift as a parameter, can be computationally expensive for large data sets.
4.1 Application to Pantheon Data
We apply our method to fit the $\Lambda$CDM and the $w$CDM models to the magnitude–redshift data of Type Ia supernovae in the Pantheon sample [24]. This data set contains 1048 spectroscopically confirmed SNIa. For all cases, we run 12 independent MCMC chains for $10^{6}$ steps. We use the Gelman-Rubin criterion of convergence. For the $\Lambda$CDM model, we achieve $R-1<0.001$, while for the $w$CDM model it is $R-1<0.005$. For $z$-errors, we use uniform priors for $v_{pi}$ in the range: $-1000\leq v_{pi}\leq 1000\,\rm km\,s^{-1}$ around the measured redshift. To verify convergence and test its rate, we increase the range of peculiar velocities to $-3000\leq v_{pi}\leq 3000\,\rm km\,s^{-1}$. Increasing this interval reduces the acceptance rate and increases $R-1$ for the fixed number of MCMC steps. However, $R-1$ is always less than $0.01$. Thus, it leads to slower convergence, but our results remain unchanged; we also obtain the same results by assuming a Gaussian distribution for peculiar velocities, imposing only the condition that $x^{*}>0$.
Parameters Priors Best-fit values
@@ 4.1 Application to Pantheon Data
$\Omega_{\rm M}$ $\mathcal{U}(0.1,0.6)$ $0.348\pm 0.035$ $0.335\pm 0.043$ $0.330\pm 0.046$
$H_{0}$ $\mathcal{U}(60,80)$ $74.07\pm 0.368$ $73.80\pm 0.546$ $73.58\pm 0.703$
$w$ $\mathcal{U}(-2,0)$ $-1.24\pm 0.14$ $-1.17\pm 0.17$ $-1.14\pm 0.18$
@@
It should be noted that our results are insensitive to regularization (Eq. (2.11)). The regularization term only changes the acceptance rate, indicating that the inclusion of redshift errors does not increase the dimensionality of our fitted physical model. This is not surprising if we note that overfitting requires the fitting function to be able to pass through most of the noisy data points. This would require the fitting function to have the flexibility that Eq. (3.2) does not have, even if we convert the redshift to a variable. The function remains a monotonic function of redshift and thus does not cause overfitting of the data.
Figures 1 and 2 display the posterior distribution and the confidence regions $1\sigma$ and $2\sigma$ for the $\Lambda$CDM and $w$CDM models for the Pantheon data. Different colors represent the three different estimators used. The red color represents the standard case, $\mathcal{E}_{1}$, blue represents the linearized estimator, $\mathcal{E}_{2}$, and the green represents the MCMC estimator, $\mathcal{E}_{3}$. In both cosmological models, we have assumed uniform priors for the cosmological parameters, which are listed in Tables 1 and 2. Perusal of Figures 1 & 2 indicates that accounting for peculiar velocities in the current data does not significantly change the estimated parameters. However, note that in Table 2 for the $w$CDM model, the marginalized estimate for the dark energy parameter $w$ is found to be marginally more consistent with the $\Lambda$CDM model, indicating that even at the current level of precision, the data could indicate non-trivial dark energy if systematic errors in redshift are not corrected. As the number of supernovae increases, this correction is likely to become even more significant. Next, we consider simulated data to investigate this issue further.
Figure 3: Constraints on the $\Lambda$CDM parameters from synthetic data with magnitude error $\Delta m=0.2$ (left) and $\Delta m=0.05$ (right) The colour scheme of Figure 1 is followed.
Models Parameters Priors True values Best-fit values
[…]
$\Lambda$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.30$ $0.217\pm 0.014$ $0.240\pm 0.016$ $0.280\pm 0.020$
$H_{0}$ $\mathcal{U}(60,80)$ $73.0$ $75.60\pm 0.338$ $74.88\pm 0.391$ $73.45\pm 0.482$
$w$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.30$ $0.368\pm 0.023$ $0.356\pm 0.028$ $0.320\pm 0.048$
$H_{0}$ $\mathcal{U}(60,80)$ $73.0$ $78.01\pm 0.642$ $77.09\pm 0.849$ $74.29\pm 1.024$
$w$ $\mathcal{U}(-3,0)$ $-1.0$ $-1.89\pm 0.213$ $-1.68\pm 0.236$ $-1.22\pm 0.228$
Table 3: Priors and best-fit values for the $\Lambda$CDM and the $w$CDM parameters obtained from synthetic data with $\Delta m=0.2$.
4.2 Application to Synthetic Data
For simulations, we first consider the parameters of the Pantheon sample: 1048 supernovae in the redshift range $z\in[0.01,2.26]$ with magnitude errors, $\Delta m\simeq 0.2$ (for simulation, we assume Gaussian random errors, $\mathcal{N}(0,0.20)$). We also consider a future SNIa project with errors four times smaller in magnitude, $\mathcal{N}(0,0.05)$. This is possible with a sample 16 times larger than the Pantheon sample with the magnitude errors obtained from binning the data in redshift (this might be achieved with Zwicky Transient Facility (ZTF)[30] and the Large Synoptic Sky Survey Telescope (LSST) [31, 32], albeit with higher supernovae density at $z<0.5$, see, e.g. [33] for details). The cosmological and other parameters are fixed to the following values: $\Omega_{\rm M}=0.30$, $H_{0}=73~{}\rm km~{}s^{-1}~{}Mpc^{-1}$, $w=-1.0$, and the absolute magnitude of supernova, $M=-19.24$. To account for peculiar velocities, each supernova is assigned an error on its redshift, $\mathcal{U}(0,0.003)$, drawn from a uniform distribution.
Our main results are shown in Figures 3 and 4 together with Tables 3 and 4. A comparison of the analysis of synthetic data with Figures 1 and 2 enables us to discern the following: (a) there is bias in the estimation of cosmological parameters using estimators ${\mathcal{E}}_{1}$ and ${\mathcal{E}}_{2}$. The bias is larger if the peculiar velocity contribution is totally neglected (${\mathcal{E}}_{1}$) and is partly alleviated if the linear approximation (Appendix A) is used and (b) while the overlapping confidence regions for $\Delta m=0.2$ do not allow us to determine this bias unambiguously using the current data, the analysis of the future data without taking into account peculiar velocities of host galaxies will result in significant bias in the determination of cosmological parameters. In particular, the analysis of SNIa data without accounting for peculiar velocities could bias the nature of dark energy, an important issue from the point of view of both cosmology and theoretical physics, and much debated in recent times following DESI results (e.g. [4]).
Figure 4: Constraints on the $w$CDM parameters from synthetic data with $\Delta m=0.2$ (left) and $\Delta m=0.05$ (right).
Models Parameters Priors True values Best-fit values
[…]
$\Lambda$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.2,0.4)$ $0.30$ $0.233\pm 0.005$ $0.280\pm 0.005$ $0.298\pm 0.005$
$H_{0}$ $\mathcal{U}(60,80)$ $73.04$ $75.42\pm 0.19$ $74.03\pm 0.129$ $73.22\pm 0.118$
$w$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.2,0.4)$ $0.30$ $0.38\pm 0.010$ $0.322\pm 0.012$ $0.310\pm 0.015$
$H_{0}$ $\mathcal{U}(60,80)$ $73.04$ $77.74\pm 0.252$ $74.86\pm 0.304$ $73.41\pm 0.243$
$w$ $\mathcal{U}(-3,0)$ $-1.0$ $-1.87\pm 0.08$ $-1.20\pm 0.07$ $-1.05\pm 0.06$
Table 4: Priors and best-fit values for the $\Lambda$CDM and the $w$CDM parameters obtained from synthetic data with $\Delta m=0.05$.
5 Conclusion and Discussion

### v2

@@ II.1.1 Estimator for nonlinear models
Regularization is commonly used to fit complex models with a large number of parameters as it acts as a constraint and facilitates stable maximization of the posterior distribution (e.g. Sivia (1996)). However, as we note later, the addition or omission of this term may not always change the dimensionality of the fit, as is further explained in the context of the specific cosmological application of this method presented in the paper. It is clear that the method advocated here is more expensive to implement numerically since each independent variable is a parameter. We chose cosmological data from SNIa as a proof of concept for our proposed method.
III Magnitude–Redshift Relation
@@
Although the measurement error in the observed redshift of a supernova host galaxy is negligibly small, due to the possible peculiar motion of the host galaxy relative to a comoving observer, its observed redshift is shifted with respect to the theoretical cosmological redshift that encodes the expansion history of the universe. To estimate this shift, we note that
$\frac{\nu_{0}}{\nu_{\rm obs}}=\frac{\nu_{0}}{\nu_{\rm com}}\frac{\nu_{\rm com}}{\nu_{\rm obs}}\approx(1+{v_{p}}/{c})(1+z)$ (14)
@@
where $\nu_{\rm obs},\nu_{0}$, and $\nu_{\rm com}$ are the supernova photon frequencies observed on earth, in the rest frame of the galaxy, and by a comoving observer at the position of the galaxy; and $v_{p}$ is the component of velocity of a galaxy with respect to the comoving observer, parallel to the line of sight from the observer to the supernova host galaxy.
@@
Peculiar motions also cause beaming of light towards the observer, alter the observed flux, and cause additional k-correction (e.g. see Ellis and Baldwin (1984) for the combined impact of all these effects on the number count). The first two effects act in concert, as an object moving towards an observer would appear closer and less redshifted. The standard analysis already takes into account measurement errors in the luminosity of an SNIa, and these errors are much larger than what we expect from the peculiar motion of host galaxies. However, this is not true for redshift measurement errors that, as already noted, are small compared to peculiar motions. At low redshifts, this effect is enhanced as peculiar redshift/blueshift can become comparable to the redshift of the host galaxy. As more and more supernovae are observed, the effective statistical error in luminosity distance reduces at any redshift because of statistical averaging. Thus, we need techniques that allow us to handle errors on redshift, even if they are small.
@@
$m=M+5\log\left(\frac{D_{\rm L}}{\text{Mpc}}\right)+25\;.$ (15)
@@
$D_{\rm L}=\frac{c(1+z)}{H_{0}}\int_{0}^{z}\frac{dz^{\prime}}{h(z^{\prime})},$ (16)
@@
$h(z)=\left[\Omega_{\rm M}(1+z)^{3}+\Omega_{\rm DE}(1+z)^{3(1+w)}\right]^{1/2}\,,$ (17)
where $\Omega_{\rm M}$ and $\Omega_{\rm DE}$ are the present-day matter and dark energy densities, respectively. The dark energy and pressure are related through the parameter $w$ through $p_{\rm DE}=w\rho_{\rm DE}$. If dark energy is assumed to be the cosmological constant, then $w=-1$ in this parametrization. The apparent magnitude can be written as
$m=M$ $+$ $5\log(cH^{-1}_{0}/{\rm Mpc})$ (18)
$+$ $5\log\left[(1+z)\int_{0}^{z}\frac{dz^{\prime}}{h(z)}\right]+25\,.$
@@
IV Results
In cosmological studies, our primary concern is with the parameters $w$ and $\Omega_{\rm M}$. Although distant SNIa are not ideal for determining $H_{0}$ due to degeneracy with $M$, we also present results for the Hubble parameter for the $\Lambda$CDM model. To break the degeneracy, we use the absolute magnitude $M=-19.24$ corresponding to the SH0ES Riess et al. (2022) value of $H_{0}=73.04$ km/s/Mpc for the Pantheon sample.
In our analysis, we consider three distinct estimators to present our results.
  • The first estimator labeled $\mathcal{E}_{1}$ is based on the standard method that does not account for redshift errors and treats them as perfectly measured.
  • The second estimator labeled $\mathcal{E}_{2}$ is based on the approximate linearized magnitude–redshift relation that includes the redshift error using the likelihood function (Eq. (24)) derived in the Appendix A. This estimator assumes that the redshift errors are Gaussian.
  • The MCMC-based final estimator $\mathcal{E}_{3}$ uses Eq. (13) described in Section II.1.1, which includes the contribution of peculiar velocity to the observed redshift by treating each redshift as a uniformly distributed random number in a range around the measured value of the redshift.
Note that only the line-of-sight component of the peculiar velocity contributes to the change in redshift. By the isotropy of the peculiar velocity dispersion, we have $\sigma_{v_{\rm 3D}}=\sqrt{3}\sigma_{v_{p}}$. Thus, any assumption we make about the magnitude of peculiar velocity must keep this in mind. For estimator $\mathcal{E}_{3}$, the distribution $P_{X}$ is minimally assumed to be uniform in a specified redshift range. More specifically, we assume that each observed $z_{i}$ in the data deviates from its true value $z^{*}_{i}$, which falls within a specified uncertainty range $z^{*}_{i}\in[z_{i}^{\rm min},z_{i}^{\rm max}]$, where the min and max values are chosen to correspond to our assumptions about peculiar velocities. We also need to assign a value to $\sigma_{z_{i}}$ in Eq. (24). We do that by setting the range chosen for $\mathcal{E}_{3}$ to be three times $\sigma_{z_{i}}$. This ensures that the one-sigma range from $-\sigma_{z_{i}}$ to $\sigma_{z_{i}}$ for a Gaussian distribution containing 99.7% probability is equal to the redshift range allowed in $\mathcal{E}_{3}$. Therefore, $\sigma_{z_{i}}=(z_{i}^{\rm max}-z_{i}^{\rm min})/6$
For presenting our results, we need to marginalize over various parameters, which means integrating over them in the posterior distribution. Also note that in Eq. (13), we need to integrate over the variable $x^{*}$, which corresponds to the redshift $z$ in this case. Clearly, it is not feasible to numerically integrate so many variables. Therefore, to carry out marginalization, we employ the MCMC sampling technique. Specifically, we develop and implement the Metropolis-Hastings algorithm within the MCMC framework to draw samples from the posterior distribution and use GetDist Lewis (2019) python library for the analysis of MCMC samples.
We found that MCMC sampling requires a longer time to converge if we treat cosmological parameters and redshift parameters on par. However, the efficiency of convergence is significantly improved by drawing multiple samples of redshifts for the same cosmological parameters. Thus, in all the results below, MCMC samples the redshift five times more often than the cosmological parameters. This ensures that MCMC is able to optimally sample the parameter space.
Figure 1: The figure displays one-dimensional posterior distributions and two-dimensional (1- and 2-$\sigma$) contours of $\Lambda$CDM parameters from Pantheon data using three different estimators. Estimator $\mathcal{E}_{1}$ (red) ignores any contribution to redshift from the peculiar motion of supernovae host galaxies, $\mathcal{E}_{2}$ (blue) considers it for the linearized model (see Appendix A for details), while $\mathcal{E}_{3}$ (green) represents our estimator, which incorporates the contribution from peculiar motion for the exact model. The errors on the estimated parameters are larger for estimators $\mathcal{E}_{2}$ and $\mathcal{E}_{3}$.
@@ IV Results
${\mathcal{E}_{1}}$ ${\mathcal{E}_{2}}$ ${\mathcal{E}_{3}}$
$\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.285\pm 0.013$ $0.296\pm 0.015$ $0.306\pm 0.016$
$H_{0}$ [km/s/Mpc] $\mathcal{U}(60,80)$ $73.86\pm 0.23$ $73.59\pm 0.30$ $73.36\pm 0.31$
Table 1: Priors and best-fit values for the spatially-flat $\Lambda$CDM model using Pantheon data. Results are shown for three different estimators: $\mathcal{E}_{1}$ ignores peculiar velocities, $\mathcal{E}_{2}$ includes them in the linear approximation, and $\mathcal{E}_{3}$ incorporates them exactly.
IV.1 Application to Pantheon Data
We apply our method to fit the $\Lambda$CDM and the $w$CDM models to the magnitude–redshift data of SNIa in the Pantheon sample Scolnic et al. (2018). This data set contains 1048 spectroscopically confirmed SNIa.
Using $\Delta z=({v_{p}/c})(1+z)$, the standard deviation used in $\mathcal{E}_{2}$ is given by $\sigma_{z}=(\sigma_{v_{p}}/c)(1+z)$. Throughout, we use $\sigma_{v_{p}}=1000\;\mathrm{km/s}$. For the estimator $\mathcal{E}_{3}$, a uniform prior $[-3000,3000]\;\rm km/s$ has been used for the peculiar velocity $v_{p}$.
For all cases, we run 12 independent MCMC chains for $10^{5}$ steps. We use the Gelman-Rubin criterion of convergence. For estimators $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$, we achieve convergence with $R-1<0.01$, while for $\mathcal{E}_{3}$ we obtain $R-1<0.1$. It might be surprising that we are able to find the optimal solution even with more than a thousand parameters. To achieve this, we need to start our MCMC runs from the observed values of redshifts. However, as long as the redshifts are assumed to have peculiar velocity components not exceeding 3000 km/s, the starting point does not matter too much. The extension of this interval reduces the acceptance rate of MCMC and increases $R-1$ for a fixed number of steps. Thus, it leads to slower convergence. Since the cosmological line-of-sight peculiar velocities are expected to be much smaller than 3000 km/s, we have not investigated convergence by relaxing our assumptions. Moreover, as noted before, for faster convergence, we split the parameter space into cosmological and redshift spaces and sample the redshift space five times more frequently than the cosmological parameter space.
We find that our results are not altered significantly by regularization (Eq. (13)). The regularization term only changes the acceptance rate and marginally changes the final $\chi^{2}$ per degree of freedom, indicating that the inclusion of redshift errors does not increase the dimensionality of our fitted physical model. This is not surprising if we note that overfitting requires the fitting function to be able to pass through most of the noisy data points that are displaced vertically due to errors in magnitude. This would require the fitting function to have the flexibility that Eq. (15) does not have with respect to treating the redshift as a parameter. The function remains monotonic, ensuring that the convergence is robust even without regularization.
Figures 1 and 2 display the posterior distribution and the confidence regions $1\sigma$ and $2\sigma$ for the $\Lambda$CDM and $w$CDM models for the Pantheon data. Estimators are denoted by different colors. The red color represents the standard case, $\mathcal{E}_{1}$, blue represents the linear approximation estimator, $\mathcal{E}_{2}$, and green represents the MCMC estimator, $\mathcal{E}_{3}$. In both cosmological models, we have assumed uniform priors for the cosmological parameters, which are listed in Tables 1 and 2. Perusal of Figures 1 and 2 indicates that accounting for peculiar velocities in the current data does not significantly change the estimated parameters. Its main impact is a small increase in errors in the estimated parameters. We further note that in Table 2 for the $w$CDM model, the marginalized estimate for the dark energy parameter $w$ is found to be marginally more consistent with the $\Lambda$CDM model, indicating that even at the current level of precision, the data could indicate evolving dark energy if systematic errors in redshift are not corrected. As the number of supernovae increases, this correction is likely to become even more significant. Next, we consider simulated data to investigate this issue further.
@@ IV.1 Application to Pantheon Data
${\mathcal{E}_{1}}$ ${\mathcal{E}_{2}}$ ${\mathcal{E}_{3}}$
$\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.350^{+0.037}_{-0.029}$ $0.330^{+0.051}_{-0.033}$ $0.320^{+0.057}_{-0.034}$
$H_{0}$ [km/s/Mpc] $\mathcal{U}(60,80)$ $74.34\pm 0.37$ $74.03\pm 0.58$ $73.83\pm 0.74$
$w$ $\mathcal{U}(-2,0)$ $-1.24^{+0.15}_{-0.13}$ $-1.16^{+0.18}_{-0.16}$ $-1.11^{+0.20}_{-0.17}$
Table 2: Priors and best-fit values for the spatially-flat $w$CDM model using Pantheon data. The results are shown for three different estimators: $\mathcal{E}_{1}$ ignores peculiar velocities, $\mathcal{E}_{2}$ includes them in the linear approximation, and $\mathcal{E}_{3}$ incorporates them exactly.
IV.2 Application to Synthetic Data
Figure 3: Constraints on the $\Lambda$CDM parameters from synthetic data with magnitude error $\sigma_{m}=0.2$ (left) and $\sigma_{m}=0.05$ (right) The colour scheme of Figure 1 is followed. As in Figure 2, the errors on the estimated parameters are marginally larger for estimators $\mathcal{E}_{2}$ and $\mathcal{E}_{3}$. In addition, the analysis of synthetic data shows that the neglect of peculiar velocities might bias the estimation of cosmological parameters. The Figure in conjunction with Figure 1 shows that this bias might not play an important role in the current data (left panel), but it would become important in the interpretation of the future data (right panel).
@@ IV.2 Application to Synthetic Data
${\mathcal{E}_{1}}$ ${\mathcal{E}_{2}}$ ${\mathcal{E}_{3}}$
$\Lambda$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.30$ $0.325\pm 0.018$ $0.313\pm 0.020$ $0.306^{+0.020}_{-0.022}$
$H_{0}$ [km/s/Mpc] $\mathcal{U}(60,80)$ $73.0$ $72.04\pm 0.34$ $72.34\pm 0.41$ $72.55\pm 0.46$
$w$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.30$ $0.287^{+0.067}_{-0.049}$ $0.317^{+0.065}_{-0.042}$ $0.331^{+0.061}_{-0.040}$
$H_{0}$ [km/s/Mpc] $\mathcal{U}(60,80)$ $73.0$ $71.85\pm 0.46$ $72.48\pm 0.68$ $72.80^{+0.73}_{-0.97}$
$w$ $\mathcal{U}(-2,0)$ $-1.0$ $-0.92^{+0.16}_{-0.12}$ $-1.05^{+0.21}_{-0.17}$ $-1.13^{+0.25}_{-0.18}$
Table 3: Priors and best-fit values for the $\Lambda$CDM and $w$CDM parameters obtained from synthetic data with $\sigma_{m}=0.2$. Results are shown for three estimators: $\mathcal{E}_{1}$, which ignores peculiar velocities; $\mathcal{E}_{2}$, which uses the linear approximation; and $\mathcal{E}_{3}$, which models them exactly.
We evaluated the efficacy of our estimators by testing them on synthetic data. In order to compare our results for the Pantheon data in the previous section with our results for synthetic data, we adopted the parameters of the Pantheon sample comprising 1048 supernovae and their redshifts that span the redshift range $z\in[0.01,2.26]$. For the first case, we fixed the quality of the data at roughly the same level as Pantheon data, assuming the supernova magnitudes to have Gaussian errors with mean zero and standard deviation $\sigma_{m}=0.2$. These results can be directly compared with the Pantheon results to determine if they exhibit similar patterns. We also consider a future SNIa survey with smaller magnitude errors $\sigma_{m}=0.05$. Such small errors are possible with a sample 16 times larger [^5] This data quality might be achieved with the Zwicky Transient Facility (ZTF) Graham et al. (2019) and the Large Synoptic Sky Survey Telescope (LSST) Abell et al. (2009); Ivezić et al. (2019), albeit with higher supernova density at $z<0.5$, see, e.g.Garcia et al. (2020) for details). than the Pantheon sample with the magnitude errors obtained by binning the data in redshift. Note that we used the Pantheon redshifts even for this case to allow for easy comparison.
Each supernova in this redshift range is assigned a peculiar velocity that is assumed to be a Gaussian random number with zero mean and standard deviation $500\,\rm km\,sec^{-1}$. As mentioned before, in order to quantify the effect of reduced errors on magnitude alone, we used the same redshifts as used in the first case. Moreover, the random errors used to generate the data are also the same, except for a reduction of a factor of four to achieve $\sigma_{m}=0.05$. The cosmological and other parameters are fixed to the following values: $\Omega_{\rm M}=0.30$, $H_{0}=73\leavevmode\nobreak\ \rm km\leavevmode\nobreak\ s^{-1}\leavevmode\nobreak\ Mpc^{-1}$, $w=-1.0$, and the absolute magnitude of the supernova $M=-19.24$.
Our main results are shown in Figures 3 and 4 together with Tables 3 and 4. We carried out many more simulations for our investigations and the results presented here are representative of the observed trends. A comparison of our analysis of synthetic data with the input cosmology enables us to discern the following. 1) There is a random error in the estimation of cosmological parameters using the estimator ${\mathcal{E}}_{1}$. The random error changes when synthetic data are drawn with a different set of random errors, and it can either overestimate or underestimate the cosmological parameters. The figures show a typical case, but the errors for real data could be larger or smaller. 2) The estimators ${\mathcal{E}}_{2}$ and ${\mathcal{E}}_{3}$ differ only slightly and are statistically consistent with each other. Their agreement with the input values of the cosmological parameters is excellent. The linear approximation estimator ${\mathcal{E}}_{2}$ seems adequate for both current and future data. 3) The discrepancies between ${\mathcal{E}}_{1}$ and the other two estimators, ${\mathcal{E}}_{2}$ and ${\mathcal{E}}_{3}$, become more pronounced with the quality of the data as the magnitude errors of the supernovae decrease from $\sigma_{m}=0.2$ to $\sigma_{m}=0.05$. 4) The errors in the cosmological parameters increase with the inclusion of redshift errors. This is not unexpected since, for variable redshifts, a wider range of cosmological parameters can fit the data. 5) Compared with the results for $\sigma_{m}=0.2$, the Pantheon results in Figures 1 and 2 show consistent trends. This indicates that our assumption for the presence of noise due to peculiar motion of host galaxies in the Pantheon data is reasonable.
In summary, although the results from the Pantheon data do not significantly improve by taking peculiar velocities into account, we find that analysis of the future data without taking into account peculiar velocities of host galaxies may result in significant bias in the determination of cosmological parameters. In particular, the analysis of SNIa data without accounting for peculiar velocities could bias the nature of dark energy, which is an important issue from the point of view of both cosmology and theoretical physics, and has been much debated in recent times following DESI results (e.g. Abdul Karim et al. (2025)).
Figure 4: Constraints on the $w$CDM parameters from synthetic data with $\sigma_{m}=0.2$ (left) and $\sigma_{m}=0.05$ (right). The colour scheme of Figure 1 is followed. As in Figure 1, the errors on the estimated parameters are marginally larger for estimators $\mathcal{E}_{2}$ and $\mathcal{E}_{3}$. In addition, the analysis of synthetic data shows that the neglect of peculiar velocities might bias the estimation of cosmological parameters. The Figure in conjunction with Figure 2 shows that this bias might not play an important role in the current data (left panel), but it would become important in the interpretation of the future data (right panel).
@@
${\mathcal{E}_{1}}$ ${\mathcal{E}_{2}}$ ${\mathcal{E}_{3}}$
$\Lambda$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.30$ $0.314\pm 0.004$ $0.299\pm 0.006$ $0.301^{+0.005}_{-0.006}$
$H_{0}$ [km/s/Mpc] $\mathcal{U}(60,80)$ $73.0$ $72.57\pm 0.09$ $72.92\pm 0.15$ $72.88^{+0.16}_{-0.15}$
$w$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.30$ $0.269^{+0.017}_{-0.015}$ $0.320^{+0.018}_{-0.015}$ $0.323^{+0.020}_{-0.016}$
$H_{0}$ [km/s/Mpc] $\mathcal{U}(60,80)$ $73.0$ $72.31\pm 0.12$ $73.29\pm 0.33$ $73.25^{+0.41}_{-0.46}$
$w$ $\mathcal{U}(-2,0)$ $-1.0$ $-0.88\pm 0.04$ $-1.09\pm 0.07$ $-1.10^{+0.10}_{-0.08}$
Table 4: Priors and best-fit values for the $\Lambda$CDM and $w$CDM parameters obtained from synthetic data with $\sigma_{m}=0.05$. Results are shown for three estimators: $\mathcal{E}_{1}$, which neglects peculiar velocities; $\mathcal{E}_{2}$, which uses a linear approximation; and $\mathcal{E}_{3}$, which models them exactly.
V Conclusion and Discussion